OpenStudy (anonymous):

A bag contains 8 black marbles, 16 white marbles, 14 red marbles, and 10 green marbles. A marble is to be drawn from the bag without looking. What is the probability that the marble drawn will not be white? A. 1/3 B. 2/3 C. 3/4 D. 15/16

zepdrix (zepdrix):

So we've got a couple of steps here... ~We need to know how many marbles there are in `total`. ~Then .. we can determine the probability that a red marble WILL be drawn, and subtract that from 1 to get the probability that a red marble will not be drawn. If that second step is confusing, we can instead add up all of the marbles that are `not red` and divide by the total number of marbles, that will give us our probability that the marble drawn is not red.

zepdrix (zepdrix):

How many marbles in total? :)

zepdrix (zepdrix):

Oh it says white now? :3 did you change it or am i crazy? lol

OpenStudy (anonymous):

I think the answer will be A 1/3 I'm correct?

zepdrix (zepdrix):

Hmmmmm, I don't think that's correct :o Non-White marbles: 8+14+10 = 32 Total marbles: 8+16+14+10 = 48 Probability of drawing a non-white marble: 32/48 Hmm I don't think that simplifies to 1/3!

OpenStudy (anonymous):

ok make sense lol, thanks anyway

OpenStudy (anonymous):

true, but can simplifies 32/48 and get 4/6 = 2/3 so the answer will be B 2/3